# Fading in the Flow: Suppression of cold gas growth in expanding galactic outflows

### v2

@@ 2.3 Cloud identification
Previous uniform cloud-crushing simulations identify clouds as a collection of gas cells that are cold and dense. Since the cells are roughly isobaric and the background pressure is fixed in this case, a denser cell is also cooler by the same factor. This is not the case in a CC85 wind where the background pressure decreases with distance. Thus, as the cloud moves downstream in the wind, it is close to the temperature floor, but its density reduces because it is isobaric relative to the falling background pressure. At some point, the density contrast relative to the background can become smaller than a chosen density contrast. We use two thresholds for cloud identification: (i) cold grid cells with temperature $<2T_{\rm cl}=8\times 10^{4}$ K, and (ii) cold ($<2T_{\rm cl}$) and overdense ($\rho>10\rho_{\rm wind}$) cells. Criterion (ii) excludes cold gas that is not sufficiently dense relative to the background, and thus the cold+dense cloud mass decreases at late times for all runs since a locally isobaric cloud continuously reduces its density relative to the background wind (c.f. Fig. 6).
@@ 2.4 Cloud tracking scheme
Figure 3: Illustration of our novel ‘cloud-tracking’ algorithm. This algorithm is generic and can take into account a spatially varying background profile in any coordinate system. However, particular to our work, we use spherical coordinates which matches with the natural geometry of a CC85 wind. We track the center of mass of the cloud material, marked by a passive tracer. Whenever this moves by more than $1$ cell length and the leftmost position of the tracer is more than $16\ R_{\rm cl}$ from the leftmost edge of the domain, we ‘effectively’ drop the cells from the left (head), introduce an identical number of cells to the right (tail) and fill them with the CC85 solution at their respective position. This procedure dynamically clips the simulation domain from the left (wind entry) and extends it to the right (wind exit) ensuring that a relatively small computational domain can follow a moving cloud and prevents its tail to quickly leak out of the simulation box.
@@
Cloud-crushing setups generally result in an advection of the initial seed cloud along the direction of the wind. This usually causes a part of the cloud or the entirety of it to move out of the simulation domain quickly. Large simulation domains can solve this issue at the cost of computations becoming prohibitively expensive. In order to overcome this, cloud crushing simulations generally employ a continuous Galilean frame transformation, which loosely follows the center of mass of the cloud material as described in Kanjilal et al. 2021; Gronke & Oh 2018; Dutta & Sharma 2019 and Shin et al. 2008.444This procedure can also be simulated with equivalent source terms in the Euler equations and can, in principle, be extended with a time-dependent boundary condition to take into account a background having spatial variations as described and implemented by Tan et al. 2023 in Cartesian coordinates. The frame transformation procedure used in previous works changes the absolute values of the coordinate positions in the simulation domain. In Cartesian coordinates, after any frame transformation, the geometry and sizes of the grid cells, as well as the underlying fluid equations, remain unchanged, as they are independent of the absolute values of the grid positions. However, for the case of spherical coordinates (used in this work), an additional complication is that the cell size and the geometry dependent source terms in the fluid equations depend on the absolute values of the coordinate positions. Therefore, a Galilean frame transformation scheme cannot simply be implemented without a continuous change in the underlying grid. This brings additional complications to the numerical setup. On the other hand, the choice of spherical coordinates is particularly convenient to study outflows, as the grid aligns with the geometry of the background field and does not introduce numerical artifacts that cause deviation from the steady-state CC85 solution that we encountered on trying to initialize a CC85 wind in cartesian boxes.
@@
Bearing these in mind, we have developed a new algorithm of ‘cloud tracking’ that can be applied to outflows or any other generic wind-tunnel simulations with spatially varying wind condition in any coordinate system and does not involve continuously changing the underlying grid. In the remainder of this subsection, we describe the implementation of this cloud tracking scheme. A tracer field (which passively advects with the flow) tags the ‘initial’ cloud material in our numerical setup. At each time step in our simulation, we calculate the leftmost (innermost/minimum) radial extent of the passive tracer by considering cells having tracer amounts above a threshold of $10^{-4}$. Our tracking scheme tries to keep this position of the minimum radial extent of the passive tracer within a fixed distance from the leftmost (innermost) radial edge of the simulation domain. This tracking procedure, therefore, switches on or off depending on the distance between the innermost radial extent of the passive tracer and the left edge of the simulation domain, which we fix to $16\ R_{\rm cl}$.
@@
When tracking is active, we calculate the radial distance moved by the center of mass of the tracer material. Whenever this distance exceeds a pre-set integer number of cells in the radial direction (we term this as the ‘shifting threshold’ and set it to 1), we shift the fluid profile by the same number of cells radially inward. This is achieved by eliminating cells at the innermost edge and introducing identical numbers of cells down the wind at the outermost edge of the simulation domain. These newly introduced cells at the right edge of the simulation domain are assigned the background CC85 wind condition according to their distance from the center of the wind. We also update the coordinate positions and cell sizes of these newly introduced cells (in spherical geometry). Therefore, our tracking algorithm effectively clips the simulation domain, on-the-fly, from the left (wind entry) and extends the domain to the right (wind exit) as the cloud gradually travels downstream. Although, in principle, this algorithm can work for any arbitrary cell size, our numerical implementation of this tracking scheme assumes uniform grid spacing both along the radial and the angular (polar and azimuthal) directions for simplicity.
@@
This cloud tracking scheme ensures that even with relatively small sized simulation domain, most of the cloud material remains within the domain for the maximum amount of time. In addition, it takes care of the spatially changing background wind in spherical coordinates. Our cloud tracking algorithm is schematically illustrated in Fig. 3. Carried by the wind blowing from from left to right, the cloud moves to the right. In response, our scheme shifts the fluid fields to the left. It is worth noting that the number of ghost cells needed for this tracking scheme to work must be more than what is required by the reconstruction stencil, plus the ‘shifting threshold’. This ensures that the ghost cells communicated across multiple processors (for internal boundaries) contain correct values of the fluid fields required by the reconstruction stencil after any shifting by the tracking algorithm. Once the shifting is complete, one or more ghost cells at the edge will have spurious values of the fluid fields. However, as long as these cells are not utilized in the reconstruction stencil, it is safe to allow this. It is worth noting that, in contrast to the previously used cloud tracking algorithms, we never switch the frame of reference here, and the simulation is always carried out in the initial lab frame.
@@
$\chi$ $\mathcal{M}_{\rm ini}$ $T_{\rm cl}$ [K] $t_{\rm cool,mix}/t_{\rm cc}$ $d_{\rm ini}/R_{\rm cl}$ $\chi$ $\mathcal{M}_{\rm ini}$ $T_{\rm cl}$ [K] $t_{\rm cool,mix}/t_{\rm cc}$ $d_{\rm ini}/R_{\rm cl}$
@@
  • The fiducial parameters
@@ 3.1 Radiative cooling and threshold cloud size in a uniform wind
Using our simulations (in spherical coordinates), we carry out ‘vanilla’ cloud-crushing at different cloud sizes to numerically evaluate the dimensionless parameter $\delta$ in Eq. 1.1. The choice $T_{\rm mix}=\sqrt{\chi}T_{\rm cl}$ is a simple approximation (Begelman & Fabian 1990) rather than a rigorous criterion. Because of this, $\delta$ would usually depend on the detailed nature of the cooling curve and the cloud/floor temperature in a simulation. Therefore, we carry out these ‘vanilla’ cloud-crushing simulations to get an estimate of the cloud radius above which clouds grow due to strong radiative cooling in a plane-parallel wind. These simulations also serve as a set of control setups against which cloud-crushing in an expanding wind can be benchmarked. Our simulation result on ‘vanilla’ cloud-crushing which we use to numerically evaluate $\delta$ is illustrated in Fig. 15 in Appendix A. We find that $\delta\approx 6$ is where the turnover from destruction to growth regime of clouds occurs in uniform cloud crushing.
@@ 3.2 Radiative cloud-crushing in an expanding background
Figure 4: The temperature slices in the $z=0$ plane from our fiducial cloud-crushing simulation in a CC85 wind. The simulation parameters are $(\mathcal{M},\chi,t_{\rm cool,mix}/t_{\rm cc}|_{\rm ini})=(1.496,100,0.2)$; cf. Table 1. We show only a few selected snapshots that clearly demonstrate the formation of elongated cold structures starting from a spherical cloud (not shown here). Similar elongated structures of cold gas with cometary tails embedded in galactic winds have been recently observed in the M82 galaxy (Bolatto et al. 2024; Fisher et al. 2025). Unlike in a uniform wind, the cloud encounters different wind conditions along its tail as it moves downstream. The cold gas also expands orthogonal to the wind direction (see the attached scale). Follow the discussion in Section 4.1.1 & Fig. 9 for a quantitative estimate on cloud expansion. Use this link here to see a video of the full evolution (also see the evolution of the corresponding density slice following this link). Moving beyond qualitative visualization, in Fig. 6, we make a detailed quantitative evaluation of the evolution of clouds in a CC85 wind.
@@
Figure 5: The temporal evolution of cold gas mass (in units of solar mass in the left panel and initial cloud mass in the right panel) in our cloud-crushing simulations for clouds of different sizes initialized in a CC85 wind (solid lines) compared with identical simulations of cloud-crushing in a uniform wind (‘vanilla’ cloud-crushing in dashed lines of the same color as their sold counterparts). We find that the growth of cold mass is significantly suppressed in an expanding background compared to ‘vanilla’ cloud-crushing. The background wind properties and the initial position of the cloud with respect to the wind remain the same in all these simulations. In this analysis, a cell is identified as being part of the cold cloud if they have a temperature less than $2T_{\rm cl}=8\times 10^{4}\ {\rm K}$. The set of constant parameters chosen for these simulations is $(\mathcal{M},\chi)=(1.496,100)$ (cf. Table 1). Each line of different color corresponds to different $t_{\rm cool,mix}/t_{\rm cc}|_{\rm ini}\propto R_{\rm thres}/R_{\rm cl}$, which translates to different cloud sizes. The time is in Myr in the left panel and normalized by the cloud crushing time $t_{\rm cc,ini}=\sqrt{\chi}R_{\rm cl}/v_{\rm wind}$ evaluated at the initial position of the cloud (the right panel). The region of the simulation domain chosen for our analysis has wind temperatures greater than $9\times 10^{4}\ {\rm K}$, higher than the temperature below which we consider gas to be cold in our analysis. For the largest clouds ($t_{\rm cool,mix}/t_{\rm cc}|_{\rm ini}=0.1,0.2$), we truncate the evolution early after a significant portion of the initial tracer marking the cloud starts leaving our analysis domain. The largest clouds exhibit the greatest deviation in their cold mass content from their ‘vanilla’ cloud-crushing counterparts. The smallest cloud from our simulation shown in this figure ($t_{\rm cool,mix}/t_{\rm cc}|_{\rm ini}=8$) gets destroyed in both the ‘vanilla’ cloud-crushing setup and cloud-crushing in a CC85 wind. It is relatively unaffected by an expanding wind and experiences a similar evolution in its cold mass content.
@@
We select clouds that are found to be growing due to radiative cooling in a uniform wind (‘vanilla’ cloud crushing) and initialize them in an expanding CC85 wind. We fix the initial cloud position and wind properties and change the cloud size by varying $t_{\rm cool,mix}/t_{\rm cc}|_{\rm ini}$ to systematically explore the growth/destruction of clouds in an expanding background. Because of this, the two independent parameters in these cloud-crushing simulations, namely $d_{\rm ini}/R_{\rm cl}$ and $t_{\rm cool,mix}/t_{\rm cc}|_{\rm ini}$, no longer remain independent (cf. Section 2.2.3). For our analysis, we consider only the sub-domain where the background wind temperature is greater than $9\times 10^{4}\ \rm K$ and classify all gas cells in this region with temperatures less than $2T_{\rm cl}=8\times 10^{4}\ \rm K$ as cold cloud material (cf. Section 2.3). This sub-domain is also used in the visualization in Fig. 4, where we show the evolution of the temperature field on a slice in the $x-y$ plane passing through the center of our simulation domain (along the $z$-axis) using a few selected snapshots. The gradient in the background wind temperature is consistent with the adiabatic expansion of a typical CC85 wind. Similar to Gronke & Oh 2020, for cooling dominated clouds, we find that their evolution in the wind can be broadly distinguished into two different stages – (i) initial ‘tail-formation’ stage when a large relative velocity between the cloud and the wind causes the initial spherical blob of cold gas to develop a long, sheared cometary tail in the direction of the wind followed by (ii) an ‘entrained’ stage when the cold gas grows due to a gradual inflow through radiative cooling around the cloud perhaps as a local cylindrical cooling flow (Dutta et al. 2022). The scales provided with the slices in Fig. 4 also indicate the gradual expansion that the cloud has undergone perpendicular to the radial direction as it has moved down the expanding wind. We find that the cold gas is nearly isothermal, and most cloud cells have a temperature approximately equal to the floor temperature in our simulation, i.e., $4\times 10^{4}\ \rm K$. Despite having some similarities with clouds moving in a uniform wind, an expanding wind causes the cloud properties to be qualitatively different from those in a uniform wind, which we discuss further.
@@
Fig. 5 illustrates the evolution of the cold mass of clouds with different sizes seeded in an expanding wind, and also compares it with the evolution of cold mass for the same clouds in a uniform wind (‘vanilla’ cloud-crushing) with identical initial conditions. We find that the growth of cold clouds is suppressed in an expanding wind compared to a uniform wind. An important result from our simulations is that such a suppression of cold mass for growing clouds occurs across clouds of all sizes. The smallest clouds that are in the destruction regime due to shear-induced mixing in the uniform wind (and inefficient radiative cooling) exhibit a similar reduction of their cold mass content over time and are relatively unaffected by the expanding background. Unlike clouds moving in a uniform wind, these clouds moving in an adiabatically expanding wind fail to remain dense with contrasting features from the background as they travel down the wind (cf. Section 3.3 & 3.4). The reduction in the density of the cold cloud as it moves in the wind plays a crucial role in suppressing the growth of the cold mass in an expanding background. When the time in our simulation is normalized by the initial cloud-crushing time $t_{\rm cc,ini}$ (which is longer for larger clouds), the largest clouds travel further downstream compared to the smaller ones within the same number of cloud-crushing times. In other words, the change in the background wind condition for every cloud radius moved by a cloud increases with cloud size $R_{\rm cl}$. This results in the highest suppression in mass growth and departures from ‘vanilla’ cloud-crushing for the larger clouds (smaller $t_{\rm cool,mix}/t_{\rm cc}|_{\rm ini}$) in our simulations. We further investigate these effects with quantitative estimates and speculate upon their implications for observations.

### v3

@@ 2.3 Cloud identification
Previous uniform cloud-crushing simulations identify clouds as a collection of gas cells that are cold and dense. In these simulations, cold gas is found to be nearly isobaric with the hot wind and this pressure remains unchanged with time. Therefore, a denser cell is also cooler by the same factor. This is not the case in a CC85 wind, where the background pressure decreases with distance. Thus, as the cloud moves downstream in the wind, it remains at a temperature close to the temperature floor, but its density reduces because it is isobaric with background pressure that reduces downstream. At some point, the density contrast relative to the background can become smaller than a chosen density contrast. Considering this, we adopt two thresholds (that needs to be simultaneously satisfied) for cloud identification: (i) cold grid cells with temperature $2T_{\rm cl}=8\times 10^{4}$ K, and (ii) locally over-dense, i.e., $\rho>10\rho_{\rm wind}(d)$, where $\rho_{\rm wind}$ is the analytic solution of the CC85 density profile at a radial distance $d$ from the wind centre of a cell in the simulation domain. The criterion (ii) excludes cold gas that is not sufficiently dense relative to the background. Discussed elaborately in Section 3.2, we find that the cold+dense ‘contrasting’ cloud mass decreases at late times for all our simulation parameters, since a locally isobaric cloud continuously reduces its density relative to the background wind (cf. Fig. 7).
@@ 2.4 Cloud tracking scheme
Figure 3: Illustration of our novel ‘cloud-tracking’ algorithm. This algorithm is generic and can take into account a spatially varying background profile in any coordinate system. However, particular to our work, we use spherical coordinates which matches with the natural geometry of a CC85 wind. We track the centre of mass of the cloud material, marked by a passive tracer. Whenever this moves by more than 11 cell length and the leftmost position of the tracer is more than $16\ R_{\rm cl}$ from the leftmost edge of the domain, we ‘effectively’ drop the cells from the left (head), introduce an identical number of cells to the right (tail) and fill them with the CC85 solution at their respective position. This procedure dynamically clips the simulation domain from the left (wind entry) and extends it to the right (wind exit) ensuring that a relatively small computational domain can follow a moving cloud and prevents its tail to quickly leak out of the simulation box.
@@
Cloud-crushing setups generally result in an advection of the initial seed cloud along the direction of the wind. This usually causes a part of the cloud or the entirety of it to move out of the simulation domain quickly. Large simulation domains can solve this issue at the cost of computations becoming prohibitively expensive. In order to overcome this, cloud-crushing simulations generally employ a continuous Galilean frame transformation, which loosely follows the centre of mass of the cloud material as described in Kanjilal et al. 2021; Gronke & Oh 2018; Dutta & Sharma 2019 and Shin et al. 2008.444This procedure can also be simulated with equivalent source terms in the Euler equations and can, in principle, be extended with a time-dependent boundary condition to take into account a background having spatial variations as described and implemented by Tan et al. 2023 in Cartesian coordinates. The frame transformation procedure used in previous works changes the absolute values of the coordinate positions in the simulation domain. In Cartesian coordinates, after any frame transformation, the geometry and sizes of the grid cells, as well as the underlying fluid equations, remain unchanged, as they are independent of the absolute values of the grid positions. However, for the case of spherical coordinates (used in this work), an additional complication is that the cell size and the geometry dependent source terms in the fluid equations depend on the absolute values of the coordinate positions. Therefore, a Galilean frame transformation scheme cannot simply be implemented without a continuous change in the underlying grid. This brings additional complications to the numerical setup. On the other hand, the choice of spherical coordinates is particularly convenient to study outflows, as the grid aligns with the geometry of the background field and does not introduce numerical artifacts that cause deviation from the steady-state CC85 solution that we encountered on trying to initialise a CC85 wind in Cartesian boxes.
@@
Bearing these in mind, we have developed a new algorithm of ‘cloud tracking’ that can be applied to outflows or any other generic wind-tunnel simulations with spatially varying wind conditions in any coordinate system and does not involve continuously changing the underlying grid. In the remainder of this subsection, we describe the implementation of this cloud tracking scheme. A tracer field (which passively advects with the flow) tags the ‘initial’ cloud material in our numerical setup. At each time step in our simulation, we calculate the leftmost (innermost/minimum) radial extent of the passive tracer by considering cells having tracer amounts above a threshold of $10^{-4}$. Our tracking scheme tries to keep this position of the minimum radial extent of the passive tracer within a fixed distance from the leftmost (innermost) radial edge of the simulation domain. This tracking procedure, therefore, switches on or off depending on the distance between the innermost radial extent of the passive tracer and the left edge of the simulation domain, which we fix to $16\ R_{\rm cl}$.
@@
When tracking is active, we calculate the radial distance moved by the centre of mass of the tracer material. Whenever this distance exceeds a pre-set integer number of cells in the radial direction (we term this as the ‘shifting threshold’ and set it to 1), we shift the fluid profile by the same number of cells radially inward. This is achieved by eliminating cells at the innermost edge and introducing identical numbers of cells down the wind at the outermost edge of the simulation domain. These newly introduced cells at the right edge of the simulation domain are assigned the background CC85 wind condition according to their distance from the centre of the wind. We also update the coordinate positions and cell sizes of these newly introduced cells (in spherical geometry). Therefore, our tracking algorithm effectively clips the simulation domain, on-the-fly, from the left (wind entry) and extends the domain to the right (wind exit) as the cloud gradually travels downstream. Although, in principle, this algorithm can work for arbitrary cell sizes, our numerical implementation of this tracking scheme assumes uniform grid spacing both along the radial and the angular (polar and azimuthal) directions for simplicity. We start with a simulation domain of size between $400-600R_{\rm cl}$ in the radial direction, and ensure that the total tracer material marking the initial cloud never falls below $95\%$ of the initial value over our time window of analysis. This ensures that almost the entirety of the cloud from its head to its tail remains confined within the simulation domain (see the tracer evolution following the link these videos at base resolution and at one level higher). In addition, Fig. 18 shows that the cloud extent along the wind saturates which ensures that the tail of the cloud is well inside the simulation domain.
@@
This cloud tracking scheme ensures that even with a relatively small sized simulation domain, most of the cloud material remains within the domain for the maximum amount of time. In addition, it takes care of the spatially changing background wind in spherical coordinates. Our cloud tracking algorithm is schematically illustrated in Fig. 3. Carried by the wind blowing from left to right, the cloud moves to the right. In response, our scheme shifts the fluid fields to the left. It is worth noting that the number of ghost cells needed for this tracking scheme to work must be more than what is required by the reconstruction stencil, plus the ‘shifting threshold’. This ensures that the ghost cells communicated across multiple processors (for internal boundaries) contain correct values of the fluid fields required by the reconstruction stencil after any shifting by the tracking algorithm. Once the shifting is complete, one or more ghost cells at the edge will have spurious values of the fluid fields. However, as long as these cells are not utilised in the reconstruction stencil, it is safe to allow this. It is worth noting that, in contrast to the previously used cloud tracking algorithms, we never switch the frame of reference here, and the simulation is always carried out in the initial lab frame.
@@
$\chi$ $\mathcal{M}_{\rm ini}$ $T_{\rm cl}$ [K] $t_{\rm cool,mix}/t_{\rm cc}|_{\rm ini}$ $d_{\rm ini}/R_{\rm cl}$ $\chi$ $\mathcal{M}_{\rm ini}$ $T_{\rm cl}$ [K] $t_{\rm cool,mix}/t_{\rm cc}|_{\rm ini}$ $d_{\rm ini}/R_{\rm cl}$
@@
  • The mass and energy injection rates, injection radius, and pressure at the initial position of the cloud are degenerate parameters and provided to give a point of reference to the readers. Our results are broadly applicable. As long as $(\chi,\mathcal{M}_{\rm ini},T_{\rm cl},t_{\rm cool,mix}/t_{\rm cc}|_{\rm ini},d_{\rm ini}/R_{\rm cl})$ is the same, the simulations, by construction, will evolve identically.
  • The fiducial parameter set
@@
  • The base resolution for all the simulations is $R_{\rm cl}/d_{\rm cell}|_{\rm ini}=8$ at the initial position of the cloud.
@@ 3.1 Radiative cooling and threshold cloud size in a uniform wind
Using our simulations (in spherical coordinates), we carry out ‘vanilla’ cloud-crushing at different cloud sizes to numerically evaluate the dimensionless parameter $\delta$ in Eq. 1.1. The choice $T_{\rm mix}=\sqrt{\chi}T_{\rm cl}$ is a simple approximation (Begelman & Fabian 1990) rather than a rigorous criterion. Because of this, $\delta$ would usually depend on the detailed nature of the cooling curve and the cloud/floor temperature in a simulation. Therefore, we carry out these ‘vanilla’ cloud-crushing simulations to get an estimate of the cloud radius above which clouds grow due to strong radiative cooling in a plane-parallel wind. These simulations also serve as a set of control setups against which cloud-crushing in an expanding wind can be benchmarked. Our simulation result on ‘vanilla’ cloud-crushing which we use to numerically evaluate $\delta$ is illustrated in Fig. 16 in Appendix A. We find that $\delta\approx 6$ is where the turnover from the destruction regime to the growth regime of clouds occurs in uniform cloud-crushing.
@@ 3.2 Radiative cloud-crushing in an expanding background
Figure 4: The temperature slices in the $z=0$ plane from our fiducial cloud-crushing simulation in a CC85 wind. The simulation parameters are $(\mathcal{M},\chi,t_{\rm cool,mix}/t_{\rm cc}|_{\rm ini})=(1.496,100,0.2)$; cf. Table 1. We show only a few selected snapshots that clearly demonstrate the formation of elongated cold structures starting from a spherical cloud (not shown here). Similar elongated structures of cold gas with cometary tails embedded in galactic winds have been recently observed in the M82 galaxy (Bolatto et al. 2024; Fisher et al. 2025). Unlike in a uniform wind, the cloud encounters different wind conditions along its tail as it moves downstream. The cold gas also expands orthogonal to the wind direction (see the attached scale). Follow the discussion in Section 4.1.1 & Fig. 10 for a quantitative estimate on cloud expansion. Use this link here to see a video of the full evolution (also see the evolution of the corresponding density slice following this link). Moving beyond qualitative visualisation, in Fig. 7, we make a detailed quantitative evaluation of the evolution of clouds in a CC85 wind. It is worth noting that we clip the slices in our visualisation to a smaller radial extent than in the simulation for clarity.
@@
Figure 5: The temporal evolution of cold gas mass (in units of solar mass in the left panel and initial cloud mass in the right panel) in our cloud-crushing simulations for clouds of different sizes initialised in a CC85 wind (solid lines) compared with identical simulations of cloud-crushing in a uniform wind (‘vanilla’ cloud-crushing in dashed lines of the same colour as their solid counterparts). We find that the growth of cold mass is significantly suppressed in an expanding background compared to ‘vanilla’ cloud-crushing. The background wind properties and the initial position of the cloud with respect to the wind remain the same in all these simulations. In this analysis, a cell is identified as being part of the cold cloud if they have a temperature less than $2T_{\rm cl}=8\times 10^{4}\ {\rm K}$. The set of constant parameters chosen for these simulations is $(\mathcal{M},\chi)=(1.496,100)$ (cf. Table 1). Each line of different colour corresponds to different $t_{\rm cool,mix}/t_{\rm cc}|_{\rm ini}\propto R_{\rm thres}/R_{\rm cl}$, which translates to different cloud sizes. The time is in Myr in the left panel and normalised by the cloud-crushing time $t_{\rm cc,ini}=\sqrt{\chi}R_{\rm cl}/v_{\rm wind}$ evaluated at the initial position of the cloud (the right panel). The region of the simulation domain chosen for our analysis has wind temperatures greater than $9\times 10^{4}\ {\rm K}$, higher than the temperature below which we consider gas to be cold in our analysis. For the largest clouds ($t_{\rm cool,mix}/t_{\rm cc}|_{\rm ini}=0.1,0.2$), we truncate the evolution early after a significant portion of the initial tracer marking the cloud starts leaving our analysis domain. The largest clouds exhibit the greatest deviation in their cold mass content from their ‘vanilla’ cloud-crushing counterparts. The smallest cloud from our simulation shown in this figure ($t_{\rm cool,mix}/t_{\rm cc}|_{\rm ini}=8$) gets destroyed in both the ‘vanilla’ cloud-crushing setup and cloud-crushing in a CC85 wind. It is relatively unaffected by an expanding wind and experiences a similar evolution in its cold mass content.
Figure 6: The temporal evolution of the centre of mass velocity of cold gas relative to the wind velocity at the position of the centre of mass of the cloud in the wind (in units of the wind velocity) in our cloud-crushing simulations for clouds of different sizes initialised in a CC85 wind. In this analysis, a cell is identified as being part of the cold cloud using the same definition as in Fig. 5 and the simulation sub-domain used in the analysis is also the same as in Fig. 5. By about $30\ t_{\rm cc,ini}$, most of the clouds have have doubled in their mass content as indicated by the diamond marks along each of the curves and they have reached $\sim 70-80$ percent of the local wind velocity indicating that they have got sufficiently entrained in the wind. The evolution of clouds pre- and post-entrainment is significantly different as discussed later in Section 4.1.1.
@@
We select clouds that are found to be growing due to radiative cooling in a uniform wind (‘vanilla’ cloud-crushing) and initialise them in an expanding CC85 wind. For our analysis, we consider only the sub-domain where the background wind temperature is greater than $9\times 10^{4}\ \rm K$ and classify all gas cells in this region with temperatures less than $2T_{\rm cl}=8\times 10^{4}\ \rm K$ as cold cloud material (cf. Section 2.3). This sub-domain is also used in the visualisation in Fig. 4, where we show the evolution of the temperature field on a slice in the xyx-y plane passing through the centre of our simulation domain (along the $z$-axis) using a few selected snapshots. The gradient in the background wind temperature is consistent with the adiabatic expansion of a typical CC85 wind. Similar to Gronke & Oh 2020, for cooling dominated clouds, we find that their evolution in the wind can be broadly distinguished into two different stages – (i) initial ‘tail-formation’ stage when a large relative velocity between the cloud and the wind causes the initial spherical blob of cold gas to develop a long, sheared cometary tail in the direction of the wind followed by (ii) an ‘entrained’ stage when the cold gas grows due to a gradual inflow through radiative cooling around the cloud perhaps as a local cylindrical cooling flow (Dutta et al. 2022). The scales provided with the slices in Fig. 4 also indicate the gradual expansion that the cloud has undergone perpendicular to the radial direction as it has moved down the expanding wind. We find that the cold gas is nearly isothermal, and most cloud cells have a temperature approximately equal to the floor temperature in our simulation, i.e., $4\times 10^{4}\ \rm K$. Despite having some similarities with clouds moving in a uniform wind, an expanding wind causes the cloud properties to be qualitatively different from those in a uniform wind, which we discuss further.
@@
Fig. 5 illustrates the evolution of the cold mass of clouds with different sizes seeded in an expanding wind, and also compares it with the evolution of cold mass for the same clouds in a uniform wind (‘vanilla’ cloud-crushing) with identical initial conditions. We find that the growth of cold clouds is suppressed in an expanding wind compared to a uniform wind. An important result from our simulations is that such a suppression of cold mass for growing clouds occurs across clouds of all sizes. The smallest clouds that are in the destruction regime due to shear-induced mixing in the uniform wind (and inefficient radiative cooling) exhibit a similar reduction of their cold mass content over time and are relatively unaffected by the expanding background. Unlike clouds moving in a uniform wind, these clouds moving in an adiabatically expanding wind fail to remain dense, with contrasting features from the background as they travel down the wind (cf. Section 3.2 & 3.4). The reduction in the density of the cold cloud as it moves in the wind plays a crucial role in suppressing the growth of the cold mass in an expanding background. As the flow expands downstream, both cloud and wind densities decrease, resulting in a lower-density mixed gas that cools far less efficiently than it does upstream. When the time in our simulation is normalised by the initial cloud-crushing time $t_{\rm cc,ini}$ (which is longer for larger clouds), the largest clouds travel further downstream compared to the smaller ones within the same number of cloud-crushing times. In other words, the change in the background wind condition for every cloud radius moved by a cloud increases with cloud size $R_{\rm cl}$. This results in the highest suppression in mass growth and departures from ‘vanilla’ cloud-crushing for the larger clouds (smaller $t_{\rm cool,mix}/t_{\rm cc}|_{\rm ini}$) in our simulations.
In Fig. 6, we show the evolution of the centre of mass velocity of the cloud relative to the wind velocity at their instantaneous position. We find that the cloud mass increases to a few multiples of the initial cold mass for most of the growing clouds by the time they get entrained. The cloud velocity approaches about 708070-80 percent of the wind velocity by $\sim 30\ t_{\rm cc,ini}$. Once entrained, the cloud properties and its mass growth rate get significantly altered as discussed further in Section 4.1.1. It also has important implication on empirical modelling of cloud-wind (cf. Section 4.1.2). We further investigate all these effects with quantitative estimates and speculate upon their implications for observations.
Figure 7: Temporal evolution of the profile of cloud ($T<2T_{\rm cl}=8\times 10^{4}\ \rm K$) pressure (left panel), density (middle panel), and velocity (right panel). The median cloud profiles are shown in thin black bordered lines, while all the cloud material is shown by dots at times 0 (red), 15 (pink), 45 (yellow) $t/t_{\rm cc,ini}$. The cloud profiles are normalised by the wind properties at the initial location of the cloud. Distance down the wind is expressed in multiples of the initial position of the cloud (w.r.t. the wind centre), and also in kpc (top labels). The simulation chosen for this plot corresponds to our fiducial parameters $(\mathcal{M},\chi,t_{\rm cool,mix}/t_{\rm cc}|_{\rm ini})=(1.496,100,0.2)$; cf. Table 1. The black finely dashed lines in all the panels show the background wind profile, and the vertical lines mark the instantaneous position of the centre of mass of the cold cloud (with the same colour as their profile). The left panel demonstrates that the cold cloud always maintains an approximate local pressure equilibrium with the wind. This implies that the cloud density is proportional to the wind pressure ($n_{\rm cl}\propto d_{\rm cl}^{-2\gamma}$; cf. Eq. 3.1) at any location (because the cloud temperature is essentially fixed; cf. Fig. 9) indicated by the thin gray dotted line in the middle panel. The cloud median density profile approximately follows this gray line. The cloud velocity profile on the right panel shows that it gradually attains the wind velocity, with its tail becoming co-moving faster than the head. In addition, the black loosely dashed line in the middle panel shows the profile of $10$ times the local wind density for reference as this denotes the density cut-off used in our work to identify clouds that stand out as contrasting features with respect to their background (see Section 3.4 & Fig. 8 for details). Animated versions of this plot are available in the following links for the profiles of pressure, density, and velocity. Refer to Fig. 24 for the same plot at a higher resolution.